. If we wanted to bring the 694 kg (approx. 3/4 ton) probe back to Earth upon completing its mission, how much energy would be needed completely escape Mars’ gravity?
Answer #1

here,

mass of Mars, M = 6.39 * 10^23 kg

radius of Mars , r = 3.3895 * 10^6 m

mass , m = 694 kg

the orbital speed , v = sqrt(2*GM/r)

the kinetic energy needed , KE = 0.5 * m * v^2

KE = 0.5 * m * ( 2 * G * M /r)

KE = 0.5 * 694 * ( 2 * 6.67 * 10^-11 * 6.39 * 10^24 /(3.3895 * 10^6))

KE = 8.73 * 10^10 J